OpenStudy (mackenzie2013):

The area of a square garden is 200 square meters. Estimate the side length of the garden.

OpenStudy (anonymous):

Do you know a way to calculate the area of a square?

OpenStudy (mackenzie2013):

I dont.

OpenStudy (anonymous):

Area = s^2, looks familiar??

OpenStudy (anonymous):

\[\sqrt{200}=10 \sqrt{2}=14.1 \]

OpenStudy (mackenzie2013):

Okay I see, I had a blonde moment LOL